Asynchronous transmission protocol is configured to use 8 data bits, 1 start bit, 2 stop bits and 1 parity bit.

The protocol efficiency of this configuration is ____ %.
Answer #1

Protocol efficiency is the amount of information bits in the transmission divided by total number of bits.

Since total number of bits = 8(data bits) + 1(start bit) + 2(stop bit) + 1(parity bit) = 12

Information bits = 8

Protocol efficiency = 8/12 = 2/3 = 66.67%